Shifter leader was sjs Shifters were hy 1. After very brief store in morning, continuous beam from around noon. 2. Data runs 16192, 16193,16195,16196, 16197, 16198 running at end of shift. Pedestal runs 16189, 16190. 3. At start of good store found L1458A mainframe was off. Took us a little time to get all voltages back on. Did not notice until run 16195 that DVM was also on mainframe, these voltages were not reset until right before run 16195. No new detector problems found. _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Sat Jun 17 16:06:02 2006
This archive was generated by hypermail 2.1.8 : Sat Jun 17 2006 - 16:06:11 EDT